What is the buffer component ratio, (NO2-)/(HNO2) of a nitrite buffer that has a pH of 2.69. Ka of HNO2 is 7.1 x 10-4.
Answer #1

In this Henderson Hasselbalch equation is used.

NO2- is deprotonated so it is a conjugate base and HNO2 is an acid.

Convert Ka to pKa

Ka = -log(Ka)
Ka = -log(7.1*10-4)
Ka = 3.15

pH = pKa + log(base/acid)

2.20 = 3.15 + log([NO2-]/[HNO2])
-0.95= log([NO2-]/[HNO2])
Take antilog both sides
10^(-0.95) = 10^(log([NO2-]/[HNO2]))
([NO2-]/[HNO2]) = 0.112

So Buffer component ratio is 0.112